OpenStudy (anonymous):

help solve 1/2X=-1+sqrt(x-2)

OpenStudy (anonymous):

correction \[\frac{ 1 }{ 2 }x=-1 +\sqrt{x+2}\]

OpenStudy (anonymous):

What you tried yet @Compgroupmail ?

OpenStudy (anonymous):

I tried to bring it to the power by \[\frac{ 1 }{ 4 }x=1+(x+2)\] /is this correct?

OpenStudy (anonymous):

see (a+b)^2 is NOT equal to a^2 + b^2 only. \[\large{(\frac{1X}{2})^2 = (-1 + \sqrt{x-2})^2 }\] \[\large{\frac{x^2}{4} = 1 + x-2 - 2\sqrt{x-2} }\]

OpenStudy (anonymous):

Oh wait in your corrected question it is x+2 not x-2 so it will be : \[\large{\frac{x} = 1+x-2 -2\sqrt{x+2}}\]

OpenStudy (anonymous):

\[\large{\frac{x^2}{4} = 1+x-2 -2\sqrt{x+2}}\]

OpenStudy (anonymous):

ok but I still don't understand how you broke it down to that. Can we go through it step by step?

OpenStudy (anonymous):

Oh! Sorry I think I confused you, yes why not... so do you know the identity of (a+b)^2 = ?

OpenStudy (anonymous):

a^2+2ab+b^2

OpenStudy (anonymous):

Right! so in RHS (Right hand side) we have : -1 + sqrt{x+2} , correct?

OpenStudy (anonymous):

left side of equation \[(\frac{ 1 }{ 2 }x)^2= \frac{ 1 }{ 4 }x\]. Correct?

OpenStudy (anonymous):

right. That is the right hand side. What's next?

OpenStudy (anonymous):

Yes , so square the right hand side ... that is : \[\large{( \sqrt{x+2} - 1 )^2 }\] Can you square that?

OpenStudy (anonymous):

no. Not 100% how to do it. That's what's messing me up.

OpenStudy (anonymous):

Wait for 1 minute, I am opening google chrome as IE is quite slow

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

add 1 to both sides and square both sides \[(\frac{ x }{ 2 } +1)^2 = x +2\]

OpenStudy (anonymous):

\[ \frac{ x^2 }{4 } +x +1 = x+2\]

OpenStudy (anonymous):

x^2 =4

OpenStudy (anonymous):

ahh. I understand now :) THanks!